# Assortment Optimization Under History-Dependent Effects

### v2

@@ 3.2 The Final Formulation
Using the linear constraints (3), (4), and (7) and the exponential cone constraint (5) to represent constraint (Choice), we obtain an MIECP formulation of problem (HAP) as follows:
$\max_{\boldsymbol{x},\boldsymbol{y},\rho}$ $\sum_{i\in[N]}\frac{1}{T}\sum_{t\in[T]}r_{i}y_{i}^{t}$ (Conic)
[…]
@@
In this remark, we consider the case where the memory length is less than two. In this case, we are able to derive the tightest convex extension, also known as the convex envelope, of $\alpha_{i}(\cdot)$. Using the convex envelope instead of the natural continuous extension, we obtain the following tighter formulation than the one given by constraint (5):
$y_{i}^{t}\geq$ $\bigl{(}\alpha_{i}(\boldsymbol{e}_{1})-\alpha_{i}(\boldsymbol{0})\bigr{)}\cdot z_{i1}^{t}+\bigl{(}\alpha_{i}(\boldsymbol{e}_{2})-\alpha_{i}(\boldsymbol{0})\bigr{)}\cdot z_{i2}^{t}+\alpha_{i}(\boldsymbol{0})\gamma_{i}^{t}$ (8a)
$y_{i}^{t}\geq$ $\bigl{(}\alpha_{i}(\boldsymbol{1})-\alpha_{i}(\boldsymbol{e}_{2})\bigr{)}\cdot z_{i1}^{t}+\bigl{(}\alpha_{i}(\boldsymbol{1})-\alpha_{i}(\boldsymbol{e}_{1})\bigr{)}\cdot z_{i2}^{t}+\bigl{(}\alpha_{i}(\boldsymbol{e}_{1})+\alpha_{i}(\boldsymbol{e}_{2})-\alpha_{i}(\boldsymbol{1})\bigr{)}\cdot\gamma_{i}^{t},$ (8b)
@@
We conclude this section with a discussion of how our formulation relates to existing formulations. Recently, a large amount of research has focused on developing strong formulations for binary linear fractional programming problems and MNL-based assortment planning problems (Sen et al. 2018, Mehmanchi et al. 2019, Atamtürk and Gómez 2020, Kılınç-Karzan et al. 2023, He et al. 2023b, Chen et al. 2024). To apply these results in solving problem (HAP), we need to linearize the nonlinear functions appearing in denominators and numerators—specifically, the attraction value functions. A prevalent linearization technique replaces each attraction value function by its multilinear extension, called the Fourier Expansion in O’Donnell (2014), and then linearizes the resulting multilinear functions using the recursive McCormick relaxation (McCormick 1976, Khajavirad 2023). We combine this linearization strategy with binary linear fractional programming formulation techniques to obtain an exact formulation of problem (HAP), which we refer to as the multilinear-extension-based formulation. The complete derivation is provided in Appendix 11.1. Later, in Section 6.1, a computational experiment shows that our formulation (Conic) significantly outperforms the multilinear-extension-based formulation.
@@ 4 Cyclic Policies: Optimality and Computation
In this section, we consider the situation in which a firm plans assortments for a long period. We establish the asymptotic optimality of cyclic policies in Section 4.1 and formulate the cyclic policy problem as an MIECP in Section 4.2. In Section 4.3, we characterize the conditions under which the length of optimal cyclic policies is the memory length plus one. Finally, in Section 4.4, we show that such conditions lead to an MILP formulation for finding an optimal two-cyclic policy that is theoretically tighter than the conic formulation.
@@ 4.1 Optimality of Cyclic Policies
The assortment planning problem with an infinite horizon can be formulated as follows. The firm can select any assortment sequence (or policy) $\boldsymbol{x}=\{\boldsymbol{x}^{t}\}_{t\in\operatorname{\mathbb{Z}}_{+}}$ with elements in $\{0,1\}^{N}$. As before, the firm may face cross-period and cross-product constraints. Let $\mathcal{P}^{\infty}$ denote the feasible set of cross-period constraints (e.g., a product cannot be offered in two consecutive periods), and let $\mathcal{X}$ denote the feasible set of cross-product constraints (e.g., assortment cardinality constraints). Then the firm aims to find an optimal assortment sequence for the following problem:
$\sup_{\boldsymbol{x}}\biggl{\{}\lim_{T\to\infty}\frac{1}{T}\sum_{t\in[T]}\sum_{i\in[N]}r_{i}\pi_{i}^{t}(\boldsymbol{x}^{t},\dots,x^{t-M})\biggm{|}\boldsymbol{x}^{t}\in\mathcal{X}\cap\{0,1\}^{N}\;\operatorname{for}t\in\operatorname{\mathbb{Z}}_{+},\ \boldsymbol{x}\in\mathcal{P}^{\infty}\biggr{\}}.$ (Infty)
An assortment policy $\boldsymbol{x}=\{\boldsymbol{x}^{t}\}_{t\in\operatorname{\mathbb{Z}}_{+}}$ is cyclic if there exists a positive integer $L$ such that $\boldsymbol{x}^{t+L}=\boldsymbol{x}^{t}$ for all $t\in\operatorname{\mathbb{Z}}_{+}$. The smallest $L>0$ for which this holds is called the cycle length of $\boldsymbol{x}$. A cyclic policy $\boldsymbol{x}$ can be represented by a finite sequence of assortments $(\boldsymbol{x}^{1},\ldots,\boldsymbol{x}^{L})$, where $L$ is the cyclic length of $\boldsymbol{x}$. In contrast, whenever we say that a finite sequence of assortments $(\boldsymbol{x}^{1},\ldots,\boldsymbol{x}^{L})$ is cyclic, we are referring to the policy in which this finite sequence of assortments is repeated infinitely often.
@@
The attraction value of product $i$ at position $t$, which depends on the assortments of product $i$ offered over the memory length, can be specified as follows:
@@
$\max\Biggl{\{}\sum_{i\in[N]}\frac{1}{L}\sum_{t\in[L]}\frac{r_{i}x^{t}_{i}\alpha^{t}_{i}(x^{1}_{i},\ldots,x^{L}_{i})}{1+\sum_{j\in[N]}x^{t}_{j}\alpha^{t}_{j}(x^{1}_{i},\ldots,x^{L}_{i})}\Biggm{|}\boldsymbol{x}^{t}\in\mathcal{X}\cap\{0,1\}^{N}\operatorname{for}t\in[L],\ \boldsymbol{x}\in\mathcal{P}^{L}\Biggr{\}},$ (Cycle)
@@
The main result of this subsection relates problems  (Infty) and (Cycle). To establish this result, we represent (Infty) as a directed graph, which we refer to as the assortment graph. Each node of the graph denotes a list of $M$ assortments, indexed from the latest to the earliest assortments in memory. Each arc represents a memory transition driven by assortment decisions. Specifically, an arc starting at a node should end at a node that drops the earliest assortment of the starting node and adds a new assortment. Hence, each arc encodes an assortment decision in a period and its weight is defined as the single-period revenue generated by the new assortment. With such a graph, we can show our assortment planning problem (Infty) is equivalent to finding the maximum mean cycle of the assortment graph (Karp 1978), whose cycle length is denoted as $L^{*}$. Given $L^{*}$, we can use (Cycle) to compute the cyclic policy. We formalize this observation as the following theorem.
@@
A detailed proof is provided in the appendix. Formulation techniques developed for solving (HAP) can be used to obtain an MIECP formulation for (Cycle); see details in Section 4.2
4.2 Finding Optimal Cyclic Policies via a Mixed-Integer Exponential Cone Program
Formulation techniques developed for solving (HAP) can be used to obtain a mixed-integer exponential cone formulation for (Cycle). Since the main discrepancy between two models is the definition of attraction value functions, we only need to modify constraints in the formulation (Conic) that involve attraction value functions. In particular, we replace $\alpha_{i}(\cdot)$ in constraint (5) and (7) by $\alpha^{t}_{i}(\cdot)$, obtaining
$y^{t}_{i}$ $\geq\exp(\beta^{0}_{i})\cdot\gamma^{t}_{i}\cdot\exp\Biggl{(}\frac{\sum_{m\in[M]}\beta_{i}^{m}z_{im}^{\tau(m\mid t)}}{\gamma^{t}_{i}}\Biggr{)}$ (9)
$y^{t}_{i}$ $\leq\alpha^{t}_{i}(\boldsymbol{h}^{\pi}_{0})\cdot\gamma^{t}_{i}+\sum_{m\in[M]}\bigl{(}\alpha^{t}_{i}(\boldsymbol{h}^{\pi}_{m})-\alpha^{t}_{i}(\boldsymbol{h}^{\pi}_{m-1})\bigr{)}\cdot z_{im}^{\tau(m\mid t)}\qquad\operatorname{for}\pi\in\Pi.$
This replacement yields the following mixed-integer exponential cone formulation of (Cycle):
$\max_{\boldsymbol{x},\boldsymbol{y},\rho}$ $\sum_{i\in[N]}\frac{1}{L}\sum_{t\in[L]}r_{i}y_{i}^{t}$ (Cycle-Conic)
$\operatorname{s.t.}$ $\boldsymbol{x}\in\mathcal{P}^{L}\text{ and }\boldsymbol{x}^{t}\in\mathcal{X}\cap\{0,1\}^{N}$ $\operatorname{for}t\in[L]$
@@ 4.2 Finding Optimal Cyclic Policies via a Mixed-Integer Exponential Cone Program
4.3 Cycle Length Characterization Under the Non-overlapping Condition
It is intuitive that, in order to avoid the negative impact on a product’s utility if the product exhibits a satiation effect, a product should not be offered consecutively. In particular, we say that a cross-period constraint on a policy $\boldsymbol{x}$ is non-overlapping if $M+1$ adjacent assortments $(\boldsymbol{x}^{t},\boldsymbol{x}^{t+1},\dots,\boldsymbol{x}^{t+M})$ are non-overlapping, that is, for each product $i\in[N]$, $x^{t}_{i}+x^{t+1}_{i}+\dots+x^{t+M}_{i}\leq 1$ for every $t\in\operatorname{\mathbb{Z}}_{+}$. This non-overlapping condition is common in sequential assortment decisions (Liu et al. 2020, Chen et al. 2023b). Li et al. (2024) identify a similar non-overlapping condition on customer orders to design a polynomial algorithm for the e-commerce assortment planning problem. Under the non-overlapping condition, any cyclic policy with length $L\leq M$ must have empty sets in some periods. We do not consider such cyclic policies. For cyclic policies with length $L>M$, we find that this condition allows us to relate the length of an optimal cycle policy for solving (Infty) to the customers’ memory length as follows.
@@ 4.3 Cycle Length Characterization Under the Non-overlapping Condition
Assume that $M<N$. An $(M+1)$-cycle policy is optimal for problem (Infty) if the cross-period constraint is non-overlapping.
The condition $M<N$ guarantees that there at least exists a non-overlapping assortment planning with each assortment non-empty. Without the non-overlapping condition, the length of the optimal cyclic policy varies as the parameters change. In particular, a slight fluctuation in the input parameters, like the base utility of a product, could change the length of the optimal cycle in the absence of the non-overlapping condition. In Appendix 10.4, we provide an example to illustrate this situation.
The base utility and history-dependent effect are estimated from real data and lie in confidence intervals. Hence, given the fact that the optimal cycle length is sensitive to input parameters, finding a cycle length with robust performance may be more critical than characterizing the optimal cycle length in practice. A positive result is that although the $(M+1)$-cyclic policy is not necessarily optimal for problem  (Infty) without the non-overlapping constraint, it still provides a good approximation for solving our problem, as illustrated in the following example.
@@
In this example, we numerically test the performance of the $(M+1)$-cyclic policy. We fix the number of products $N=30$ and vary the length of the planning horizon $T\in\{10,30,50\}$. The revenue and base utility of each product are randomly generated from uniform distributions with ranges $[1,10]$ and $[-1,1]$, respectively. For a weak (W) (resp. strong (S)) satiation effect, we set the history-dependent effect $\beta_{i}^{m}$ using a uniform distribution $U[-1,0]$ (resp. $U[-2,-1]$). For each instance, we compute the revenue gap between the $(M+1)$-cyclic policy and the best feasible solution, defined as follows:
$G_{\textsf{rev,cycle}}=100\%\times\frac{R_{\mathsf{fea}}-R_{\mathsf{cycle}}}{R_{\mathsf{fea}}},$
where $R_{\mathsf{cycle}}$ is the revenue achieved by the $(M+1)$-cyclic policy and $R_{\mathsf{fea}}$ is the revenue of the best feasible solution obtained by (Conic) within $7200$ seconds.
Table 2: Revenue gap of cyclic policies
@@
The non-overlapping condition provides an additional benefit in problem formulation when cycle length is small. In the following subsection, we provide a bound-free model formulation when $M=1$ with non-overlapping requirement.
4.4 A Strong Formulation for Finding Two-Cyclic Policies
Here, we will focus on the computation of the two-cyclic policy with the non-overlapping constraint. Although the formulation (Cycle-Conic) is convenient, one can directly solve it to obtain an optimal cyclic policy. However, its computational performance depends on the bounds of no-purchase probabilities. We next propose a bound-free formulation for the two-cyclic policy, which is provably tighter than its conic counterpart (Cycle-Conic), and, in Section 6.2, we demonstrate its computational efficacy. Our formulation is built on the polyhedral structure of the relation between choice probabilities across two periods and assortment decisions.
We start by reformulating (Cycle) into a simpler form which is a special case of the assortment optimization problem studied in Caro et al. (2014). When $M=1$, attraction value functions are univariate functions over the binary domain $\{0,1\}$, and, thus, can be represented using an affine function. For each product $i\in[N]$, let $u_{i}:=\exp(\beta^{0}_{i})$ and $d_{i}:=\exp(\beta^{0}_{i}+\beta^{1}_{i})-\exp(\beta^{0}_{i})$. Then, for each period $t\in[T]$, the attraction value function can be linearized as follows:
$\alpha^{1}_{i}(x^{2}_{i})=u_{i}+d_{i}x^{2}_{i}\quad\text{ and }\quad\alpha^{2}_{i}(x^{1}_{i})=u_{i}+d_{i}x^{1}_{i}\quad\operatorname{for}\,(x^{1}_{i},x^{2}_{i})\in\{0,1\}^{2},$
where $u_{i}$ represents the base attraction value of product $i$, and $d_{i}$ is the incremental attraction value of product $i$ conditional on whether it is offered in another period. Due to the presence of the non-overlapping constraint $x_{i}^{1}+x_{i}^{2}\leq 1$, we obtain a further simplification:
$x_{i}^{t}\alpha_{i}^{t}(x_{i}^{\tau(1|t)})=u_{i}x_{i}^{t}+d_{i}x_{i}^{1}x_{i}^{2}=u_{i}x_{i}^{t}\qquad\text{ for }(x^{1}_{i},x^{2}_{i})\in\{0,1\}^{2},$
where the second equality holds since $x^{1}_{i}x_{i}^{2}=0$. Therefore, (Cycle) is equivalent to
$\max$ $\sum_{t=1,2}\frac{1}{2}\sum_{i\in[N]}\frac{r_{i}u_{i}x^{t}_{i}}{1+\sum_{j\in[N]}u_{j}x^{t}_{j}}$ (2-Cyclic)
$\operatorname{s.t.}$ $x_{i}^{1}+x_{i}^{2}\leq 1$ $\operatorname{for}i\in[N]$
$\boldsymbol{x}^{t}\in\mathcal{X}\cap\{0,1\}^{N}$ $\operatorname{for}t=1,2.$

### v3

@@ 3.1.2 The Final Formulation
Using the linear constraints (3), (4), and (7) and the exponential cone constraint (5) to represent (Choice), we obtain an MIECP formulation of problem (HAP) as follows:
$\max_{\boldsymbol{x},\boldsymbol{y},\boldsymbol{\rho},\boldsymbol{\gamma},\boldsymbol{z}}$ $\frac{1}{T}\sum_{t\in[T]}\sum_{i\in[N]}r_{i}y_{i}^{t}$ (Conic)
$\operatorname{s.t.}$ $\boldsymbol{x}^{t}\in\{0,1\}^{N}\cap\mathcal{X}\text{ and }\boldsymbol{x}\in\mathcal{P}$ $\operatorname{for}\,t\in[T]$
@@
In this remark, we consider the case where the memory length is less than two. In this case, we can derive the tightest convex extension, also known as the convex envelope, of $\alpha_{i}(\cdot)$. Using the convex envelope instead of the natural continuous extension, we obtain the following tighter formulation than the one given by constraint (5) for $M=2$. If $\boldsymbol{\beta}_{i}$ are all positive or negative, we have
$y_{i}^{t}\geq$ $\alpha_{i}(\boldsymbol{0})(\gamma_{i}^{t}-z_{i1}^{t}-z_{i2}^{t})+\alpha_{i}(\boldsymbol{e}_{1})z_{i1}^{t}+\alpha_{i}(\boldsymbol{e}_{2})z_{i2}^{t}$ (8a)
$y_{i}^{t}\geq$ $\alpha_{i}(\boldsymbol{e}_{1})(\gamma_{i}^{t}-z_{i2}^{t})+\alpha_{i}(\boldsymbol{e}_{2})(\gamma_{i}^{t}-z_{i1}^{t})+\alpha_{i}(\boldsymbol{1})(z_{i1}^{t}+z_{i2}^{t}-\gamma_{i}^{t}),$ (8b)
and if $\boldsymbol{\beta}_{i}$ are mixed, the convex envelope is
$y_{i}^{t}\geq$ $\alpha_{i}(\boldsymbol{0})(\gamma_{i}^{t}-z_{i1}^{t})+\alpha_{i}(\boldsymbol{e}_{1})(z_{i1}^{t}-z_{i2}^{t})+\alpha_{i}(\boldsymbol{1})z_{i2}^{t}$ (9a)
$y_{i}^{t}\geq$ $\alpha_{i}(\boldsymbol{0})(\gamma_{i}^{t}-z_{i2}^{t})+\alpha_{i}(\boldsymbol{e}_{2})(z_{i2}^{t}-z_{i1}^{t})+\alpha_{i}(\boldsymbol{1})z_{i1}^{t}$ (9b)
@@
We conclude this section with a discussion of how our formulation relates to existing formulations. Recently, a large amount of research has focused on developing strong formulations for binary linear fractional programming problems and MNL-based assortment planning problems (e.g. Sen et al. 2018, Mehmanchi et al. 2019, Atamtürk and Gómez 2020, Kılınç-Karzan et al. 2023, He et al. 2024, Chen et al. 2024). To apply these results in solving problem (HAP), we need to linearize the nonlinear functions appearing in denominators and numerators—specifically, the attraction value functions. A prevalent linearization technique replaces each attraction value function by its multilinear extension, called the Fourier Expansion in O’Donnell (2014), and then linearizes the resulting multilinear functions using the recursive McCormick relaxation (McCormick 1976, Khajavirad 2023). We combine this linearization strategy with binary linear fractional programming formulation techniques to obtain an exact formulation of problem (HAP), which we refer to as the multilinear-extension-based formulation. The complete derivation is provided in the ecompanion EC.5.1. Later, in Section 6.1, a computational experiment shows that our formulation (Conic) significantly outperforms the multilinear-extension-based formulation.
@@
3.2 Optimality of Revenue-Ordered Policies
A prevalent strategy for solving assortment optimization is to construct Revenue-Ordered (RO) assortments, which is an assortment including products with revenue higher than a threshold. Specifically, RO structures and their variants have been shown to be optimal for solving assortment optimization problems under the MNL choice model (Talluri and Van Ryzin 2004), sequential MNL choice models (Gao et al. 2021), MNL choice models with network effects (Wang and Wang 2017), etc. Inspired by the simplicity and optimality of RO structures, we would like to understand how it works in our problem.
We find that when there are no cross-product and cross-period constraints, if the history-dependent effect is non-negative, that is, $\boldsymbol{\beta}\geq 0$, the optimal assortment in each period is an RO assortment, and the revenue threshold increases when $T$ increases. Such an assortment planning can be obtained by a sequential revenue-ordered policy. The sequential RO policy is obtained by sequentially selecting the best RO assortment based on the currently updated product utility and maximizing revenue in the current period. This policy is adaptive while ignoring the impact of the current assortment decision on future revenue. Details of the sequential RO policy are shown in Algorithm 1. Because there are no cross-product and cross-period constraints, the computation time required to find the optimal RO assortment is $\mathcal{O}(N)$ in each period (Talluri and Van Ryzin 2004, Liu and Van Ryzin 2008). Thus, the total computation time of the sequential RO policy is $\mathcal{O}(NT)$.
Algorithm 1 Sequential revenue-ordered policy of (HAP)
1:  Input: $\boldsymbol{r},\{\beta_{i}^{0}\}_{i\in[N]},\boldsymbol{\beta},M,N,T$
2:  for $t=1,2,\dots,T$ do
5:  end for
6:  Output: $(\boldsymbol{x}^{1},\dots,\boldsymbol{x}^{T})$
Theorem 3.4
Assume the absence of cross-product and cross-period constraints. Then, the sequential-revenue-ordered policy solves (HAP) if the history-dependent effects are non-negative, that is, $\boldsymbol{\beta}\geq\boldsymbol{0}$.
For the case of negative history-dependent effects, which is the computationally intractable regime of our model, as we argued in Proposition 2.1, the sequential policy fails to solve our model. However, the RO structure still exists under the negative and even mixed history-dependent effects and appears in a different form. For an optimal assortment planning of model (HAP), we argue that the set of products that are offered across all periods has an RO structure.
Proposition 3.5
Assume the absence of cross-product and cross-period constraints. Then, there exists an optimal assortment $\boldsymbol{x}=(\boldsymbol{x}^{1},\ldots,\boldsymbol{x}^{T})$ of (HAP) such that $\bigl{\{}i\in[N]\bigm{|}\sum_{t\in T}x^{t}_{i}\geq 1\bigr{\}}$ is revenue-ordered.
@@ 4 Cyclic Policies: Optimality and Computation
In this section, we consider the situation in which a firm plans assortments for a long period. We establish the asymptotic optimality of cyclic policies in Section 4.1 and formulate the cyclic policy problem as an MIECP in Section 4.2. In Section 4.3, we characterize conditions under which the length of optimal cyclic policies is the memory length plus one. In Section 4.4, we show that such conditions lead to a tighter MILP formulation for finding optimal $(M+1)$-cyclic policies.
@@ 4.1 Optimality of Cyclic Policies
The assortment planning problem with an infinite horizon can be formulated as follows. The firm can select any assortment sequence (or policy) $\boldsymbol{x}=\{\boldsymbol{x}^{t}\}_{t\in\operatorname{\mathbb{Z}}_{+}}$ with elements in $\{0,1\}^{N}$. As before, the firm may face cross-product and cross-period constraints. Let $\mathcal{X}$ (resp. $\mathcal{P}^{\infty}$) denote the feasible set of cross-product (resp. cross-period) constraints. Then, the firm aims to find an optimal assortment sequence for the following problem:
$\sup_{\boldsymbol{x}}\biggl{\{}\lim_{T\to\infty}\frac{1}{T}\sum_{t\in[T]}\sum_{i\in[N]}r_{i}\pi_{i}^{t}(\boldsymbol{x}^{t},\dots,\boldsymbol{x}^{t-M})\biggm{|}\boldsymbol{x}^{t}\in\mathcal{X}\cap\{0,1\}^{N}\;\operatorname{for}t\in\operatorname{\mathbb{Z}}_{+},\ \boldsymbol{x}\in\mathcal{P}^{\infty}\biggr{\}}.$ (Infty)
An assortment policy $\boldsymbol{x}=\{\boldsymbol{x}^{t}\}_{t\in\operatorname{\mathbb{Z}}_{+}}$ is cyclic if there exists a positive integer $L$ such that $\boldsymbol{x}^{t+L}=\boldsymbol{x}^{t}$ for all $t\in\operatorname{\mathbb{Z}}_{+}$. The smallest $L$ for which this holds is called the cycle length of $\boldsymbol{x}$. A cyclic policy $\boldsymbol{x}$ can be represented by a finite sequence of assortments $(\boldsymbol{x}^{1},\ldots,\boldsymbol{x}^{L})$, where $L$ is the cycle length of $\boldsymbol{x}$. In contrast, whenever we say that a finite sequence of assortments $(\boldsymbol{x}^{1},\ldots,\boldsymbol{x}^{L})$ is cyclic, we are referring to the policy in which this finite sequence of assortments is repeated infinitely often.
@@
The attraction value of product $i$ at position $t$ can be specified as follows:
@@
$\max\Biggl{\{}\frac{1}{L}\sum_{t\in[L]}\sum_{i\in[N]}\frac{r_{i}x^{t}_{i}\alpha^{t}_{i}(x^{1}_{i},\ldots,x^{L}_{i})}{1+\sum_{j\in[N]}x^{t}_{j}\alpha^{t}_{j}(x^{1}_{i},\ldots,x^{L}_{i})}\Biggm{|}\boldsymbol{x}^{t}\in\mathcal{X}\cap\{0,1\}^{N}\operatorname{for}t\in[L],\ \boldsymbol{x}\in\mathcal{P}^{L}\Biggr{\}},$ (Cycle)
@@
The main result of this subsection relates to problems (Infty) and (Cycle). To establish this result, we represent (Infty) as a directed graph, which we refer to as the assortment graph. Each node of the graph denotes a list of $M$ assortments, indexed from the latest to the earliest assortments in memory. Each arc represents a memory transition driven by an assortment decision. Specifically, an arc starting at a node should end at a node that drops the earliest assortment of the starting node and adds a new assortment. Hence, each arc encodes an assortment decision in a period, and its weight is defined as the single-period revenue generated by the new assortment. With such a graph, we can show that (Infty) is equivalent to finding the maximum mean cycle of the assortment graph (Karp 1978), whose cycle length is denoted as $L^{*}$. Given $L^{*}$, we can use (Cycle) to compute the cyclic policy. We formalize this observation as the following theorem.
@@
4.2 Finding Optimal Cyclic Policies
Formulation techniques developed for solving (HAP) can be used to obtain a mixed-integer exponential cone formulation for (Cycle). Since the main discrepancy between the two models is the definition of attraction value functions, we only need to modify constraints in (Conic) that involve attraction value functions and historical assortments. In particular, we replace $x^{t-m}_{i}$ in (4) by $x_{i}^{\tau(m|t)}$ and $\alpha_{i}(\cdot)$ in (7) by $\alpha^{t}_{i}(\cdot)$:
$\max\bigl{\{}0,\ \rho^{t}_{U}x^{\tau(m|t)}_{i}+\gamma^{t}_{i}-\rho^{t}_{U}\bigr{\}}\leq z_{im}^{t}\leq\min\bigl{\{}\gamma_{i}^{t},\rho^{t}_{U}x^{\tau(m|t)}_{i}\bigr{\}}$ $\quad\operatorname{for}\,m\in[M]$ (10a)
$y^{t}_{i}\leq\alpha_{i}^{t}(\boldsymbol{w}^{\sigma}_{i,0})(\gamma^{t}_{i}-\tilde{z}_{i\sigma(1)}^{t})+\sum_{k\in[M]}\alpha_{i}^{t}(\boldsymbol{w}^{\sigma}_{i,k})(\tilde{z}_{i\sigma(k)}^{t}-\tilde{z}_{i\sigma(k+1)}^{t})$ $\quad\operatorname{for}\,\sigma\in\Omega,$ (10b)
where $\boldsymbol{w}^{\sigma}_{i,k}\in\{0,1\}^{L}$, and its $\tau(m|t)^{\text{th}}$ element equals the $m^{\text{th}}$ element of $\boldsymbol{h}_{i,k}^{\sigma}$ defined in (7) and others are zero. This replacement yields the following mixed-integer exponential cone formulation of (Cycle):
$\max_{\boldsymbol{x},\boldsymbol{y},\boldsymbol{\rho},\boldsymbol{\gamma},\boldsymbol{\boldsymbol{z}}}$ $\frac{1}{L}\sum_{t\in[L]}\sum_{i\in[N]}r_{i}y_{i}^{t}$ (Cycle-Conic)
$\operatorname{s.t.}$ $\boldsymbol{x}^{t}\in\mathcal{X}\cap\{0,1\}^{N}\text{ and }\boldsymbol{x}\in\mathcal{P}^{L}$ $\operatorname{for}t\in[L]$
@@ 4.2 Finding Optimal Cyclic Policies
4.3 Cycle Length Characterization
It is intuitive that, in order to avoid the negative impact on a product’s utility if the product exhibits a satiation effect, a product should not be offered consecutively. In particular, we say that a cross-period constraint on a policy $\boldsymbol{x}$ is non-overlapping if $M+1$ adjacent assortments $(\boldsymbol{x}^{t},\boldsymbol{x}^{t+1},\dots,\boldsymbol{x}^{t+M})$ are non-overlapping, that is, for each product $i\in[N]$, $x^{t}_{i}+x^{t+1}_{i}+\dots+x^{t+M}_{i}\leq 1$ for every $t\in\operatorname{\mathbb{Z}}_{+}$. This non-overlapping condition is common in assortment optimization literature (Liu et al. 2020, Chen et al. 2023b). For instance, Li et al. (2024) identify a similar non-overlapping condition on customer orders to plan assortments for a local warehouse. Under the non-overlapping condition, any cyclic policy with length $L\leq M$ must have empty sets in some periods. We do not consider such cyclic policies. For cyclic policies with length $L>M$, we find that this condition allows us to relate the length of an optimal cyclic policy for solving (Infty) to the customers’ memory length as follows.
@@ 4.3 Cycle Length Characterization
Assume that $M<N$. An $(M+1)$-cyclic policy is optimal for problem (Infty) if the cross-period constraint is non-overlapping.
We remark that Theorem 4.2 holds for mixed satiation-addiction effects, because its proof replies on a graph whose structure unaffected by signs of $\boldsymbol{\beta}$.
The condition $M<N$ guarantees that there at least exists a non-overlapping assortment policy without empty assortments. Without the non-overlapping condition, the optimal cycle length varies as the parameters change. In particular, a slight fluctuation in the input parameters, like the base utility of a product, could change the optimal cycle length. We provide an example to illustrate this situation in EC.4.4.
The base utility and history-dependent effect are estimated from real data and lie in confidence intervals. Hence, given the fact that the optimal cycle length is sensitive to input parameters, finding a cycle length with robust performance may be more critical than characterizing the optimal cycle length in practice. A positive result is that although the $(M+1)$-cyclic policy is not necessarily optimal for (Infty) without the non-overlapping constraint, it still provides a good approximation, as illustrated in the following example.
@@
We numerically test the performance of the $(M+1)$-cyclic policy under satiation effects. We fix the number of products $N=30$ and vary the planning horizon $T\in\{10,30,50\}$. The revenue and the base utility of each product are randomly generated from uniform distributions with ranges $[1,10]$ and $[-1,1]$, respectively. For a weak (W) (resp. strong (S)) satiation effect, we set the history-dependent effect $\beta_{i}^{m}$ using a uniform distribution $U[-1,0]$ (resp. $U[-2,-1]$). For each instance, we compute the revenue gap between the $(M+1)$-cyclic policy and the best feasible solution, defined as $100\%\times\frac{R_{\mathsf{fea}}-R_{\mathsf{cycle}}}{R_{\mathsf{fea}}}$, where $R_{\mathsf{cycle}}$ is the revenue achieved by the $(M+1)$-cyclic policy and $R_{\mathsf{fea}}$ is the revenue of the best feasible solution obtained by (Conic) within $7200$ seconds.
Table 1: Revenue gap of $(M+1)$-cyclic policies under satiation effects with $N=30$ and no constraints
@@
4.4 A Strong Formulation for Finding Non-overlapping $(M+1)$-Cyclic Policies
In this subsection, we propose a bound-free formulation for the $(M+1)$-cyclic policy, which is provably tighter than its conic counterpart (Cycle-Conic) whose computational performance depends on the bounds ($\rho^{t}_{L}$ and $\rho^{t}_{U}$) of the no-purchase probability variable $\rho^{t}$.
From the non-overlapping constraint $x_{i}^{t}+x^{\tau(1\mid t)}_{i}+\dots+x^{\tau(M\mid t)}_{i}\leq 1$, we have:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{i}^{t}\alpha_{i}^{t}(x^{1}_{i},\ldots,x^{L}_{i})=x_{i}^{t}\exp\biggl{(}\beta_{i}^{0}+\sum_{m\in[M]}\beta^{m}_{i}x^{\tau(m\mid t)}_{i}\biggr{)}=x_{i}^{t}\exp(\beta_{i}^{0})\qquad\text{ for }t\in[M+1].}$
Let $u_{i}=\exp(\beta_{i}^{0})$.  (Cycle) is equivalent to
$\max$ $\frac{1}{M+1}\sum_{t\in[M+1]}\sum_{i\in[N]}\frac{r_{i}u_{i}x^{t}_{i}}{1+\sum_{j\in[N]}u_{j}x^{t}_{j}}$ ((M+1)-Cyclic)
$\operatorname{s.t.}$ $x_{i}^{t}+\sum_{m\in[M]}x^{\tau(m\mid t)}_{i}\leq 1$ $\operatorname{for}i\in[N]$
$\boldsymbol{x}^{t}\in\mathcal{X}\cap\{0,1\}^{N}$ $\operatorname{for}t\in[M+1].$